OpenStudy (anonymous):

70% of. 279. Is what number write. An equation

OpenStudy (unklerhaukus):

\[70/100 \times 279=195.3\]

OpenStudy (anonymous):

0.7*279=195.2

OpenStudy (anonymous):

oops 195.3…wrong key ; )